Three forces with magnitudes of 68 pounds, 108 pounds, and 131 pounds act on an object at angles of 30°, 45°, and 120° respectively, with the positive x-axis. Find the direction and magnitude of the resultant force. (Round your answers to one decimal place.)

direction °
magnitude    lb
Answer #1

X component, Fx = 68 cos30 + 108 cos45 - 131 cos60 = 69.76 lb

Y component, Fy = 68 sin30 + 108 sin45 + 131 sin60 = 223.82 lb

Resultant force, F = sqrt(Fx^2 + Fy^2) = 234.44 lb

Direction, theta = arctan (Fy/Fx) = 72.7 degree

Comment in case any doubt please rate my answer ....